Have We Reached the Top Yet

Pekka K. Sinervo University of Toronto / IPP

1. The Standard Model

2. Why a Top Quark?

3. The CDF Experiment

4. Evidence from the Most Recent Search

5. What’s Next?

University of Guelph November 1994